How many moles of PCl5 can be produced from 59.0 g of Cl2 (and excess P4)?
Express your answer to three significant figures and include the appropriate units.
1
Expert's answer
2015-10-20T04:21:33-0400
The balanced chemical eqiation is
P4 + 10Cl2 = 4PCl5[sub]
The mass of clorine is m(Cl[sub]2
) = 59.0 g
The amount of clorine is n(Cl2) = m(Cl2)/M(Cl2) = 59/70.906 = 0.832 mol
The reacion stoichiometry shows that
10 moles of Cl2 yield 4 moles of PCl5;
0.832 of Cl2 yield x moles of PCl5
n(PCl5) = 4*0.832/10 = 0.333 mol

Answer: 0.333 mol